# Self-Supervised Skeleton-Based Action Representation Learning: A Benchmark and Beyond

### v2

@@ 4.5 Benchmark on Action Recognition Task
Method xview xsub
1% 10% 1% 10%
Semi-Supervised Methods
X-CAR [81] - 78.2 - 76.1
MAC-Learning [65] - 78.5 - 74.2
Self-Supervised Methods
LongT GAN [100] - - 35.2 62.0
MS2L [40] - - 33.1 65.1
ISC [72] 38.1 72.5 35.7 65.9
CMD [52] 53.0 80.2 50.6 75.4
HiCo [16] 54.8 78.3 54.4 73.0
PCM3 [98] 53.1 82.8 53.8 77.1
PCM3++ 57.5 84.5 57.1 79.4
[…]
4.6 Benchmark on Different Downstream Tasks
To give a thorough evaluation of generalization capacity of current SSL models and our method PCM3++, we conduct extensive experiments and perform benchmarking on five downstream tasks, including action recognition, action retrieval, occluded recognition, action detection, and few-shot learning. For comparison, we choose the popular SSL methods mostly based on GRU model, which achieves a strong performance currently, while also ensuring a fair comparison with our GRU-based approach.
1) Skeleton-based Action Recognition. The following two evaluation settings are adopted.
Firstly, the unsupervised learning results using linear evaluation protocol is shown in Table LABEL:tab:main_comp, and PCM3++ achieves better or competitive results compared with other state-of-the-art methods on different evaluation protocols. Meanwhile, the results on PKUMMD dataset can be found in supplementary material, which also indicates the strong generalization capacity on the noisy data.
In addition, we also report the semi-supervised setting results, which fine-tunes the whole model/classifier with only a portion of labeled data. As shown in Table 2, PCM3++ obtains better performance than both the semi- and self-supervised methods with different proportions of training data, and only fine-tuning the classifier tends to be better with extremely little data. In contrast, previous works based on reconstruction [100] or contrastive learning [72, 52] solely can not achieve satisfactory results.
Table 3: The results of action retrieval with joint stream.
Method NTU 60 (%) NTU 120 (%)
xsub xview xsub xset
ISC [72] 62.5 82.6 50.6 52.3
HiCLR [97] 60.6 73.1 46.0 46.0
CRRL [78] 60.7 75.2 - -
CMD [52] 70.6 85.4 58.3 60.9
HaLP [61] 65.8 83.6 55.8 59.0
HiCo [16] 67.9 84.4 55.9 58.7
MAMP [51] 62.0 70.0 - -
UmURL [70] 71.3 88.3 58.5 60.9
PCM3 [98] 73.7 88.8 63.1 66.8
PCM3++ 75.4 89.4 64.5 67.1
Table 4: The results of occluded action recognition.
Method Occluded NTU 60
Spatial Occ. Temporal Occ.
xsub xview xsub xview
MoCo-GRU [26] 64.8 72.6 68.8 74.8
ISC [72] 62.8 70.6 68.9 76.8
CRRL [78] 56.8 61.4 61.0 66.2
CMD [52] 67.1 72.7 72.7 79.5
HiCo [16] 66.4 72.1 71.2 76.7
PCM3 [98] 80.8 87.0 77.6 86.1
PCM3++ 81.8 88.0 79.9 85.8
2) Skeleton-based Action Retrieval. Following previous work [68], a K-nearest neighbors (KNN) classifier ($k$=1) is adopted to retrieve the nearest training sample for each testing data in the representation space. The precision is reported as accuracy in Table 3. Our method achieves a competitive performance compared with other latest methods. This indicates the highly distinguishable representations learned by our method. Meanwhile, it is found that contrastive learning based methods usually perform better than reconstruction-based methods, e.g., LongT GAN [100] and MAMP [51].
Table 5: Action detection results on PKUMMD Part I benchmark under linear evaluation protocol.
Method mAP@tIoU (%)
0.1 0.3 0.5
Supervise 63.4 61.4 53.4
Unsupervised Methods
MoCo-GRU [26] 68.2 67.2 63.4
CRRL [78] 57.6 55.7 52.1
ISC [72] 64.6 62.9 58.7
CMD [52] 73.7 72.8 68.4
HiCo [16] 51.8 50.8 45.8
LAC [88] 55.2 - -
PCM3 [98] 73.3 72.8 68.2
PCM3++ 75.5 74.6 69.8
Table 6: Comparison of few shot skeleton-based action recognition. ($k$, $n$) represents $k$-way $n$-shot.
Method xset (%) xsub (%)
(5, 1) (5, 5) (5, 1) (5, 5)
Supervise 64.5 81.2 67.3 85.2
Unsupervised Methods
MoCo-GRU [26] 58.5 78.7 61.1 81.3
CRRL [78] 58.3 77.1 55.2 75.0
ISC [72] 62.3 81.3 64.4 83.6
CMD [52] 64.9 82.8 68.6 85.7
HiCo [16] 64.6 81.7 67.1 84.2
PCM3 [98] 65.2 82.9 68.8 86.8
PCM3++ 66.7 84.8 69.9 87.3
3) Action Recognition with Occlusion. We evaluate the transfer ability of representations learned from the clean dataset to the action recognition with occluded data. We adopt the linear evaluation protocol. Following [67], we construct a synthetic occluded dataset on both spatial and temporal dimension. For spatial dimension, different body parts are randomly masked. For temporal dimension, we set a random block of frames to zeros. The testing set is constructed by the same masks across different methods, with a masking ratio of [0.3, 0.7].
As presented in Table 4, our method can capture underlying structures in the distorted data owing to the masked contrastive learning, and is good at dealing with the spatial occlusion. Meanwhile, this capacity extends well for temporal occlusion, demonstrating a desirable generalization capacity. However, due to the lack of modeling learning of occluded data, other methods are with poor robustness for occlusion.
4) Skeleton-based Action Detection. We follow the previous works [49, 43], and evaluate the short-term motion modeling capacity by action detection task. We train the attached linear classifier to predict frame-level categories to produce the final proposal. The encoder is pre-trained on NTU 60 dataset and then we fix it, transferring the learned representations to untrimmed PKUMMD Part I dataset. The mean average precision (mAP) of different actions is adopted as metric with different temporal overlapping ratios (tIoU) in Table 5. We provide a fully fine-tuned model (train both the encoder and classifier) without pre-training as the baseline. First, it can be found the SSL pre-training can significantly improve the performance compared with the baseline model. Besides, our method integrates the clip-level representation contrastive learning and achieves the best scores compared with latest CMD [52], HiCo [16], and LAC [88], which ignore the clip-level representation modeling.
@@ 4.6 Benchmark on Different Downstream Tasks
5) Unsupervised Few-shot Learning. In this task, we evaluate the performance of SSL as few-shot learners follow the previous work [48]. Specifically, the model is first pre-trained on NTU 60 dataset. Then, a simple classifier, e.g., Support Vector Machine (SVM) in our implementation, is fitted on the output features by the pre-trained encoder of the support data set. Finally, the adapted classifier along with the encoder is utilized to infer the query samples. We select 20 new categories in NTU 120111Specifically, the classes (index from 0) 60, 61, 66, 69, 72, 78, 79, 80, 84, 90, 91, 95, 96, 98, 99, 100, 102, 106, 108, 111, 113 114, 115 are selected. which are not seen in pre-training, as the support and query sets following [46]. In Table 6, our method generalizes well on the unseen categories and surpasses the supervised baseline and other methods notably, although there is no training for unseen categories.
6) Compared with the Baseline Method. We evaluate and present the comparison results of our baseline model, PCM3 [98], on various downstream tasks as shown in Table LABEL:tab:main_comp-6. As our new improvement, we integrate an effective clip-level contrastive learning scheme and present a novel post-distillation training strategy. These methodological advancements consistently improve the model’s representation quality and robustness on various downstream tasks, especially the semi-supervised action recognition and action detection performance.
To sum up, self-supervised learning provides a strong solution to boost various downstream tasks for versatile skeleton-based action understanding. Generally, a higher recognition performance roughly indicates a better representation space for most downstream tasks, as found in Table 3 and Table 6, but not equal to greater robustness to occlusion. Meanwhile, action detection is another crucial task for action analysis, which mainly focus on the short-term motion modeling. We design a novel asymmetric clip-level contrastive learning pretext task and effectively boost the detection performance. Remarkably, our proposed method achieves promising results on these tasks compared with other methods that ignore the versatile representation learning of different granularity.
@@ 4.7 Analysis of PCM3++
1) Analysis of clip contrastive learning. We add the temporal clips as the asymmetric positive sample of the anchor sequence, which further improves the representation quality as shown in Table 7. Remarkably, the clip-level contrastive learning significantly boost the action detection performance by promoting short-term modeling capacity. Our designs boost versatile representation learning of different granularity and the effect on more downstream tasks can be found in Supp.
2) Effect of the post-distillation design. As shown in Table 7, the post-distillation training strategy can bring further improvement slightly for the recognition task. It alleviates the false negative problem by removing the one-hot pseudo label in the contrastive learning, achieving a more compact representation space.
Table 8: Results of different backbones.
Method Backbone xsub xview
ActCLR [41] GCN 80.9 86.7
CSTCN [77] GRU 83.1 88.7
HiCLR [97] Transformer 76.6 80.8
PCM3++ GCN 80.6 85.1
PCM3++ GRU 84.8 91.0
PCM3++ Transformer 80.2 84.9
Table 9: FLOPs and Params results.
[…]
GL-Transformer [33] 214M 59.4G 83.8%
ISC [72] 106M 13.7G 85.2%
CMD [52] 99M 17.3G 86.9%
PCM3 [98] 103M 15.0G 90.4%
PCM3++ 103M 17.9G 91.0%
3) Results of different backbones. We give more results with different backbones in Table 8. Specifically, for GCN and Transformer, we adopt ST-GCN [86] and DSTA-Net [64] following previous works [41, 97]. As we can see, our method shows good generalization capacity across different model backbones, and performs better or on-par compared with the latest SSL methods. Additional improvement can be potentially achieved by further searching for hyper-parameters on different backbones, which is, however, not the focus of this paper, and we leave it for future work. We finally adopt GRU model in implementation for the more friendly memory usage of GPU and the higher performance.
4) Complexity Analysis. We give an analysis of space and computational complexities of our method for pre-training in Table 9. Compared with other GRU-based and Transformer-based methods, our method achieves a significant performance improvement with an acceptable cost of the complexity. For the space complexity, the main extra cost is the reconstruction decoder (4M of space). For the computational complexity, the encoding process of different positive samples contributes most of the overhead.
@@ 5 Conclusions and Future Directions
  • Multi-Modal Learning. The exploration of multi-modal pre-training including human skeleton data is still insufficient, e.g., text, infrared data, and RGB images. For example, as a useful complement to skeleton, RGB data can provide additional background knowledge that can boost action representation learning.
@@
  • Versatile Representation Learning. As discussed in previous sections, most works only focus on the recognition task, leaving the generation capacity across different tasks of SSL models under-explored. Meanwhile, new frameworks can also be explored, e.g., Diffusion model [27], which is promising to handle both generative and discriminative downstream tasks.
@@
  • Competing interests. This work was supported in part by the National Natural Science Foundation of China under Grant No.62172020, and in part by the Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology).
  • Data availability. The adopted NTU-RGB+D [62, 44] and PKU-MMD [43] dataset are popular public benchmarks in skeleton-based action understanding. The code for data processing can be found in https://github.com/JHang2020/PCM3.

### v3

@@ 4.4.3 Cross-Dataset Analytics for Different Downstream Tasks
$\bullet$ In contrast, for the tasks requiring to explicitly generalize to new data distributions, e.g., temporal occluded recognition with distribution shift by impairing data structure (Note only spatial occlusion is seen in the pre-training) and few shot tasks with new categories, introducing new categories can benefit the performance significantly. This indicates more essential knowledge of human motion is enabled for model to boost generalization.
To sum up, we analyze the effect of pre-training on different datasets to downstream tasks from two factors, category distribution and the sample number within each class, providing valuable insights for the skeleton SSL dataset choice and construction. Overall, a better performance can be expected with a larger pre-training dataset. However, it is also necessary to consider the actual downstream scenarios and the adopted pre-training methods to adjust data component.
Table 3: Ablation study on the novel positive pair and semantic guidance.
Novel Positive Pair Semantic Guidance xview
$s_{mask}$ $s_{predict}$
88.2
89.3
90.0
90.5
@@ 4.5 Analysis of PCM3++
1) Evaluating the PCM3++. As shown in the previous results in Table LABEL:tab:main_comp and Table LABEL:tab:other_tasks, the proposed method achieves remarkable results on different downstream tasks compared with the state-of-the-art methods, and demonstrates significant improvement over our baseline method [zhang2023prompted]. As our new improvement, we integrate an effective clip-level contrastive learning scheme and present a novel post-distillation training strategy. These methodological advancement is studied as follows.
2) Analysis of clip contrastive learning. We add the temporal clips as the asymmetric positive sample of the anchor sequence, which further improves the representation quality as shown in Table 4. Remarkably, the clip-level contrastive learning significantly boost the action detection performance by promoting short-term modeling capacity. Our designs boost versatile representation learning of different granularity and the effect on more downstream tasks can be found in Supp.
3) Effect of the post-distillation design. As shown in Table 4, the post-distillation training strategy can bring further improvement slightly for the recognition task. It alleviates the false negative problem by removing the one-hot pseudo label in the contrastive learning, achieving a more compact representation space.
Table 5: FLOPs, params, and training time (all tested on a single NVIDIA A40 GPU) results.
Model Params FLOPs Training Acc.
Time
GL-Trans. [kim2022global] 214M 59.4G 53.62h 83.8%
ISC [thoker2021skeleton] 106M 13.7G 45.49h 85.2%
CMD [mao2022cmd] 99M 17.3G 44.13h 86.9%
PCM3 [zhang2023prompted] 103M 15.0G 45.31h 90.4%
PCM3++ 103M 17.9G 47.95h 91.0%
4) Novel positive pair and semantic guidance. We analyze the effect of synergetic design between masked prediction and contrastive learning in Table 3. As we can see, the masked views and predicted views as positives bring notable improvement, indicating the potential benefits of masked prediction training to contrastive learning. On the other hand, we propagate the gradients of contrastive branch to masked prediction as high-level semantic guidance rather than stop the gradient. This boosts masked prediction training and yields more informative positive samples, leading to further performance improvement.
5) Complexity analysis. We give an analysis of space and computational complexities of our method for pre-training in Table 5. Compared with other GRU-based and Transformer-based methods, our method achieves a significant performance improvement with an acceptable cost of the complexity and training time. For the space complexity, the main extra cost is the reconstruction decoder (4M of space). For the computational complexity, the encoding process of different positive samples contributes most of the overhead.
@@ 5 Conclusions and Future Directions
  • Multi-Modal Learning. The exploration of multi-modal pre-training including human skeleton data is still insufficient, e.g., texts, and RGB images. For example, as a useful complement to skeleton, RGB data can provide additional background knowledge that can boost action representation learning. Some recent works [petrovich2023tmr, yu2024mopatch] explore the human motion-text representation learning. However, the insufficiency of high-quality motion-text datasets still limits the model generalization capacity.
@@
  • Versatile Representation Learning. As discussed in previous sections, most works only focus on the recognition task, leaving the generation capacity across different tasks of SSL models under-explored. Meanwhile, new frameworks can also be explored, e.g., Diffusion model [ho2020denoising], which is promising to handle both generative and discriminative downstream tasks.
@@
  • Competing interests. This work was supported in part by the Beijing Major Science and Technology Project under Contract No. Z251100008425023, in part by the National Natural Science Foundation of China under Grant 62471009, and in part by the Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology).
  • Data availability. The adopted NTU-RGB+D [shahroudy2016ntu, liu2019ntu] and PKU-MMD [liu2020pku] dataset are popular public benchmarks in skeleton-based action understanding. The code for data processing can be found in https://github.com/JHang2020/PCM3.